OpenStudy (anonymous):

Determine whether the problem should be solved using the formula for an arithmetic sequence, arithmetic series, geometric sequence, or geometric series. Explain your answer in complete sentences. You do not need to solve. Jackie deposited $5 into a checking account in February. For each month following, the deposit amount was doubled. How much money was deposited in the checking account in the month of August?

OpenStudy (amistre64):

its prolly best to just dig into it and start writing a recurrsion for it

OpenStudy (amistre64):

m1 = d m2 = d+d m3 = d+d+d+d m4 = d+d+d+d+d+d+d+d d 2d 4d 8d 16d , let d=1 for the time being 1 2 4 8 16 32 64 .... = 2^(n-1)